m / s 1 in 8 seconds of motion, starting from a body pause. Covers a distance of 2 m. What will be the acceleration of the body?
1
Expert's answer
2020-01-03T09:23:08-0500
"s=0.5at^2"

The acceleration of the body:


"a=\frac{2s}{t^2}=\frac{2(2)}{8^2}=0.0625\frac{m}{s^2}"